CONFORMATION AND DIASTEREOMERIC SPECIFIC SPECTROSCOPIC INVESTIGATION OF α/β-PEPTIDES Ac-Phe-β^{3}-hAla-NHMe AND Ac-β^{3}-hAla-Phe-NHMe

\underline{WILLIAM H. JAMES III}, ESTEBAN E. BAQUERO, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907; SOO HYUK CHOI and SAMUEL H. GELLMAN, Department of Chemistry, University of Wisconsin-Madison, Madison, WI 53706.

β-peptides differ from naturally occurring α-peptides in having an extra carbon linking amide groups in the polypeptide chain. While the conformational preferences and spectroscopic signatures of α-peptides have received much attention, far less is known about β-peptides. In collaboration with the Gellman group at UW-Madison, we have undertaken a detailed study of small α/β-peptides containing an aromatic chromophore that enables their study via UV double resonance spectroscopy. This talk will describe our progress on the single-conformation spectroscopy of α/β-peptides, focusing particular attention on Ac-Phe-β^{3}-hAla-NHMe (1) and the (L)- and (D)-Phe diastereomers of Ac-β^{3}-hAla-Phe-NHMe (2L and 2D). The molecules are brought into the gas phase by thermal heating (240 ^{∘}C) and cooled in a supersonic expansion. One color, resonant two-photon ionization spectroscopy and UV-UV holeburning are used to determine the number of conformations present and obtain their ultraviolet spectral signatures. Ultraviolet transitions due to six conformations each of (1),(2L), and (2D), respectively, have been observed. Resonant ion-dip infrared spectroscopy in the amide NH stretch region has been carried out. The amide NH stretch region is particularly diagnostic of the number and type of intramolecular H-bonds present. These are denoted by the number of atoms involved in the H-bonded ring so formed (e.g., C6 = a 6-atom H-bonded ring). Spectroscopic evidence will be presented for the presence of a variety of H-bonded architectures, including C5/C8, C5/C6, and C7/C8 double ring motifs. The comparison of these experimental results with those on small β-peptides and small ACPC ((S,S)-trans-2-aminocyclopentanecarboxylic acid) containing α/βpeptides will be made. As an additional point of comparison, DFT and MP2 calculations provide a basis of assignment to specific H-bonded families and point out the challenges to current theory to quantitatively describe the variety of amide-amide H-bonds that can be formed.